OpenStudy (anonymous):

I just don't know what I'm doing wrong, I keep getting 21.82 again! Please help?!? If an object is thrown in an upward direction from the top of a building 160 ft. high at an initial speed of 21.82 mi/h., what is its final speed when it hits the ground?(Disregard wind resistance. Round answer according to significant figures and do not reflect negative direction in your answer. a)110 b)106 c)80

OpenStudy (anonymous):

hmm.. can u show me how u keep getting that answer?

OpenStudy (anonymous):

So i keep getting 8.0 rounded to the closest tenth it should be distance/over time, but to be honest i would have mashy look at it

OpenStudy (anonymous):

alright, let me see if I can show you how I've been trying to figure it out

OpenStudy (anonymous):

ok so, scratch the 21.82 answer, I was plugging it in wrong, but is the answer 80 then?

OpenStudy (anonymous):

what is the value of g that u take?

OpenStudy (anonymous):

I used 9.8m/s2

OpenStudy (anonymous):

Im pretty sure its 8.0 but again i would wait for mashy to answer

OpenStudy (anonymous):

Its just that, that's not one of the choices. I am an online student, so I have no teacher. A lot of times, its really hard to learn physics by yourself. So its problems like this one, where I just don't know how to arrive at the answer. Im really bad with formulas because I don't always know where to plug things in. The choices on the program don't match that of which you are getting, its just really confusing

OpenStudy (anonymous):

yeah im online too, umm whats the formula your using

OpenStudy (anonymous):

The attached file is the formulas for the topic im on now. I've been trying a few of them, because I don't know which one is the proper one to use

OpenStudy (anonymous):

ill take a look one sec

OpenStudy (anonymous):

thats alot of equations ill try a few and tell you what i get your using 9.8 for g right?

OpenStudy (anonymous):

yeah because isn't g the acceleration due to gravity(9.8)

OpenStudy (anonymous):

haha.. g is 9.8 METERS PER SECOND square but ur units are in FPS.. so you need to convert it into feet per second square..

OpenStudy (anonymous):

so its 32.15 feet/ second square!

OpenStudy (anonymous):

Are we done? Use m g h' = (1/2) m V^2 to go from initial velocity to height above initial position. Use m g h = (1/2) m v^2 to go from h = h' + 160 ft to the velocity v at the ground, v.

OpenStudy (anonymous):

none of your answers are right and yet idk the right answer :(